OpenStudy (anonymous):

one ninth of sales at subway are cash, if cash sales were 690 what r total sales

OpenStudy (anonymous):

cash sales for the week are 690

OpenStudy (anonymous):

Let the total sales at subway be X, And X/9= cash sales = 690. X/9=690 => X = 9 * 690 X = 6210. Therefore the total sales is 6210

OpenStudy (anonymous):

tks